OpenStudy (spinex12):

What are the first four terms in the multiplication pattern given by the formula? 4 · 3n A. 4, 12, 36, 108 B. 7, 21, 63, 189 C. 12, 36, 108, 324 D. 12, 24, 36, 48

hartnn (hartnn):

just plug in n=1,2,3,4 in 3n what do u get ?

hartnn (hartnn):

\(4\times 3n\)

OpenStudy (spinex12):

12

hartnn (hartnn):

you mean 12n ?

OpenStudy (spinex12):

yes

hartnn (hartnn):

12*1 = ..? 12*2 = ..? 12*3 = ..? 12*4 = ..?

OpenStudy (spinex12):

ok I get it thank you :)

hartnn (hartnn):

welcome ^_^